OpenStudy (anonymous):

Please help! Find the following infinite sum, if it exists. If one does not exist enter "no infinite sum". (LOOK AT RESPONSE)

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}6\left( 2/3 \right)^{n-1}\]

OpenStudy (anonymous):

is it n^(-1) or is it n-1? :)

OpenStudy (anonymous):

n-1

hero (hero):

\[\sum_{n =1}^{\infty} 6(\frac {2}{3})^{n-1}\]

hero (hero):

What is going on with Latex? That looks like crap

OpenStudy (anonymous):

Yes Hero, that is correct. My bad :/

OpenStudy (anonymous):

alright, I think I got it :\[\sum_{n=1}^{\infty} 6(\frac{2}{3})^{n-1} = \sum_{n=1}^{\infty} 2^n\] then you can use the root test to proceed ^_^ got it?

OpenStudy (anonymous):

since you'll be left with :\[2(2)^{n-1} = 2^{n-1+1} = 2^n\] right? .-.

OpenStudy (anonymous):

or is that wrong? hmmm

OpenStudy (anonymous):

hero what do you think? ^_^

OpenStudy (anonymous):

I'm not sure... Hero?

hero (hero):

Let me see if I can verify that...

hero (hero):

I get 18 for the sum

OpenStudy (anonymous):

meh, then that means my steps were wrong ^_^"?

OpenStudy (anonymous):

Hmm. I'll try. Thank you for the help guys :)

OpenStudy (anonymous):

np and good luck ! ^_^

OpenStudy (anonymous):

18 is correct - I usually find it easiest to look at the first few terms, and from there figure out what the first term (a) and the common ratio (r) are. In this case we have: 6, 4, 8/3, 16/9 etc..., so a=6 and r=2/3. Then use the geometric series formula: Sum = a/(1-r) = 6/(1-2/3) = 18